Hello, I was reviewing radicals for a test tomorrow and had trouble with one question on the bonus worksheet.
It was.....
√16x^2
This would be a simple problem, but I'm dreadful at factoring and cant remember what to do with the 16x^2 or how to solve it.
Thank you so much for your time, work, and effort!!

Answer by jim_thompson5910(35256)

Since 16 = 4^2, this means that 16x^2 = 4^2*x^2 = (4x)^2


So the square root of 16x^2 is the same as the square root of (4x)^2


So the square and the square root effectively cancel out leaving just 4x


So √16x^2 = 4x where x is nonnegative


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim